OpenStudy (anonymous):

slope 3/4, y intercept: 9 write an equation of the line with the given slope and y-intercept

OpenStudy (anonymous):

if anyone wouldn't mind helping me out.

OpenStudy (anonymous):

Y intercept is when X = 0. It's another way of saying one of the points on the line is (0,9). m (slope) = (3/4) Slope intercept form is: y=mX+b, where m = slope First, solve for the Y intercept (which will get you b): 9 = m(0)+b Then substitute b and m into y=mx+b.